OpenStudy (anonymous):

find the missing numerator. 1/x-1/x+3=??/x(x+3) I will give a MEDAL!

OpenStudy (studygurl14):

\[\frac{ 1 }{ x - 1 }/{x + 3} = \frac{ ? }{ x(x + 3) }\] Is that it?

OpenStudy (anonymous):

no, the last is correct. but the first part is 1/x - 1/x+3

OpenStudy (studygurl14):

oh, sorry. then this: \[\frac{ 1 }{ x } - \frac{ 1 }{ x } + 3 = \frac{ ? }{ x(x + 3) }\] ?

OpenStudy (anonymous):

but the x+3 is in parenthesis on the denominator of the second one

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

the second one is 1/ (x+3)

OpenStudy (studygurl14):

sorry. \[\frac{ 1 }{ x } - \frac{ 1 }{ x + 3 } = \frac{ ? }{x(x + 3) }\] Okay, this equation makes much more sense. lol

OpenStudy (anonymous):

yeahhh lol

OpenStudy (studygurl14):

Okay, so first step. You have to change the fractions so that they have the same denominator. Do you know how to do that with variables?

OpenStudy (anonymous):

no, i've basically forgotten all of this, and i'm taking my final now ):

OpenStudy (anonymous):

thats why i came here lol to help me figure out how to do it

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

\[\frac{ 1(x + 3) }{ x(x + 3) } - \frac{ 1(x) }{ (x + 3)x} = \frac{ ? }{ x(x + 3) }\] Now you can add the numerators because the denominators are the same.

OpenStudy (anonymous):

ohh okay

OpenStudy (studygurl14):

can you do the rest?

OpenStudy (anonymous):

yes, its three correct?

OpenStudy (studygurl14):

yes, good! So ? = 3